Mr. Hall will use fencing to build a rectangular pen to hold his goats. He plans to use exactly 400 feet of fencing. Which function gives the area, A, of the pen based on its width, w?
How to I figure this out?

Found 2 solutions by addingup, ikleyn:
Answer by addingup(3677)

2W + 2L = 400 there are a number of combinations that will make this statement true.
Example: 2(75) + 2(125) = 150 + 250 = 400 in this example, the area would be 75*125 = 9,375

Answer by ikleyn(52799)

.
If w is the width, then the length is 200-x,   and the area is


    A(w) = width*length = w*(200-w) = -w^2 + 200w.


It is the function of w the problem asks for.

Solved.